6. Compute f(3) if f(x) = (. Vat2 +1 dt

help
Answer #1

solution : f(x)= d4 +2+, at P(a) from lebinite theoror, f(x) = Ju(222) 27, .2x - J4 (2) 27 .32² f(x) = (J 4x4 + 1) 2x - (4x +if you have any further doubts regarding this please feel free to ask.